OpenStudy (anonymous):

lim n->∞ (1-1/4)*(1-1/9)*...*(1-1/n^2)

OpenStudy (anonymous):

@quickstudent @waterineyes @Elsa213 @radar @TheAsker2002 @Yolo4mecuite

Elsa213 (elsa213):

r there any answer choices?

OpenStudy (dan815):

hmmm

OpenStudy (anonymous):

hm

Elsa213 (elsa213):

@Alexander95

OpenStudy (dan815):

0 -.-

OpenStudy (dan815):

maybeenot

OpenStudy (dan815):

1/2?

OpenStudy (anonymous):

NOT.

Elsa213 (elsa213):

cx

OpenStudy (dan815):

lol

OpenStudy (dan815):

the answer is definately les than 3/4 ill tell u that

Elsa213 (elsa213):

^^^^^^

Elsa213 (elsa213):

less*

Elsa213 (elsa213):

The*

Elsa213 (elsa213):

definitely*

Elsa213 (elsa213):

I'll*

Elsa213 (elsa213):

you*

OpenStudy (anonymous):

@Alexander95 is your keyboard working?

Elsa213 (elsa213):

that.*

OpenStudy (anonymous):

no answer choice but I found out that those expressions can be simplified

OpenStudy (dan815):

umm well

OpenStudy (dan815):

its prollyyyy 1/2 gut feeling

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }*\frac{ 8 }{ 9 }*\frac{ 15 }{ 16}*...*\frac{ n^2-1 }{ n^2 }\]

OpenStudy (dan815):

yeah

OpenStudy (anonymous):

but why 1/2 @dan815 instead of 1

OpenStudy (dan815):

itd not 1/2, its a guess, it cant be one since you keeep multiplying by a smaller fraction

OpenStudy (dan815):

hmm lemme tink i think i know a way to solve this

OpenStudy (dan815):

this cud be solve by eigen values

OpenStudy (anonymous):

dan815 you're right this product really tends to 1/2

OpenStudy (anonymous):

Base is larger is every term, so every term is less than 1, so overall will be more less result.

OpenStudy (anonymous):

I.ve multiplied the first 20 terms of this exp But still need an algebraic solvation of the problem

OpenStudy (zarkon):

\[\frac{ 1\times 3 }{ 2\times 2 }*\frac{ 2\times 4 }{ 3\times 3 }*\frac{ 3\times 5 }{ 4\times 4}*...*\frac{ (n-1)\times(n+1) }{ n\times n }\]

OpenStudy (zarkon):

then start canceling

OpenStudy (dan815):

pretty cool

OpenStudy (anonymous):

\[= \lim_{n \to \infty}\frac{1}{2} \times \frac{n+1}{n}\]

OpenStudy (anonymous):

Wow, Zarkon, Great...!! :)

OpenStudy (anonymous):

So, Dan was right.. :)

OpenStudy (dan815):

i think i found a cool soln* with eigen values too xD, if u write recurrence relation, prolly too much work for this algebra problem tho